OpenStudy (anonymous):

A pancake recipe requires one and one third cups of milk to one cup of flour. If one and one half cups of milk is used, what quantity of flour will be needed, according to the recipe?

OpenStudy (lilg132):

one and 1/6

OpenStudy (lilg132):

cups of flour

OpenStudy (lilg132):

dont forget to click best response :)

OpenStudy (anonymous):

incorrect..

OpenStudy (lilg132):

=O

OpenStudy (lilg132):

not sure then

OpenStudy (anonymous):

The milk in the new recipe is 9/8 of the amount in the original. That indicates the flour needs to increase by the same ratio, so I'd say you need about a cup and an eighth.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Original recipe: 1 1/3 cups milk = 8/6 cups; second recipe: 1 1/2 cups of milk = 9/6 cups. Ratio is 9/8, increase flour by the same ratio.